OpenStudy (vera_ewing):

What is the equation in point-slope form of the line passing through (-2, -5) and (2, 3)?

OpenStudy (vera_ewing):

@jim_thompson5910

OpenStudy (vera_ewing):

@Directrix

Directrix (directrix):

(-2, -5) and (2, 3) --> What is the slope of the line determined by these two points? @vera_ewing

OpenStudy (vera_ewing):

2

Directrix (directrix):

Yes.

OpenStudy (vera_ewing):

A. (y + 2) = -2(x + 5) B. (y - 2) = 2(x - 3) C. (y - 3) = 2(x - 2) D. (y + 3) = -2(x + 2)

OpenStudy (vera_ewing):

so it's C right?

Directrix (directrix):

point-slope form y - y1 = m(x - x1) where m is the slope of the line and )x1,y1) is a point on the line.

OpenStudy (vera_ewing):

Oh so it's B?

Directrix (directrix):

(2,3) is a point on the line y - 3 = 2* ( x -2) Is that an option.

OpenStudy (vera_ewing):

Yeah that's option c. I was right :) thank you! :)

Directrix (directrix):

You are welcome.